OpenStudy (anonymous):

Hope I didn't make any mistakes. \[\int_c cosydx+x^2sinydy\] C is the rectangle with vertices (0,0),(5,0),(5,2) and (0,2) \[p=cosy\] \[\frac{dp}{dy}=-siny\] \[q=x^2siny\] \[\frac{dq}{dx}=2xsiny\] \[\int\int_R (2xsiny+siny)dxdy\] \[\int_0^2\int_0^5(2xsiny+siny)dxdy\] \[\int_0^2[x^2siny]_0^5dy\] \[\int 25sinydy\] 25[-cos(2)+cos(0)]

OpenStudy (anonymous):

just a little mistake in integrating for x i think...

OpenStudy (anonymous):

are you using green's theorem ?? i guess .

OpenStudy (anonymous):

yes Greens theorem

OpenStudy (anonymous):

It should have been dydx?

OpenStudy (anonymous):

C is a rectangle...it does'nt matter to write dydx or dxdy

OpenStudy (anonymous):

my first integration is wrong?

OpenStudy (anonymous):

a little mistake \[\int\limits (2x \sin y+\sin y )dx = \sin y (x^2+x) +A\]

OpenStudy (anonymous):

so in the last integral for y that will be ...30 sin y ... instead of ...25 sin y...

OpenStudy (anonymous):

yes mukusha did figure out after first integration and application of limit you should have 30\[30\int\limits_{0}^{2}\sin(y)dy\]

OpenStudy (anonymous):

Oh ok. Thank you guys!

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

it's perfect. I don't see any errors in it. :X